OpenStudy (czesc):

INSTANT FAN//MEDAL Find the polar coordinates for point P which has Cartesian coordinates: (1,-sqrt(3))

jimthompson5910 (jim_thompson5910):

cartesian points are in the form (x,y) polar points are in the form (r,theta) you'll use these two formulas r = sqrt(x^2 + y^2) theta = arctan(y/x)

jimthompson5910 (jim_thompson5910):

in this case, x = 1, y = -sqrt(3)